OpenStudy (anonymous):

2+squareroot of x+5=4 extraneous soloution

OpenStudy (dls):

\[\LARGE (2+\sqrt{x+5})=4\] what about this?

OpenStudy (anonymous):

wht do u want to about it

OpenStudy (dls):

\[\LARGE \sqrt{x+5}=2\] \[\LARGE x+5=4=>x=1\]